OpenStudy (anonymous):

arkababu: if we add 1 to the numerator and subtract 1 from denominator of a fraction then the fraction becomes 1.It also become 1/2 if we only add 1 to the denominator. What is the fraction.

OpenStudy (shubhamsrg):

new numerator be n and denominator be d we are given (n+1)/(d-1) =1 and (n+1)/d = 1/2 2 eqns, 2 variables,you know what to do.. best of luck ! :P